Update 21/07/2016 17:00 The power outage has now been resolved and service has been restored as normal. Update 21/07/2016 10:10 Our supplier has said the fault has been identified as being part of the ongoing major BT outage within one of their London data centers. All traffic has been moved to unaffected carries so customers should now see service return to normal. Update 21/07/2016 09:38 An issue has been identified […]

Read More →

Update 21/07/2016 17:00 Our ISP has said that engineers have completed the required work and service should now be restored. Update 21/07/2016 12:00 Our ISP has said that engineers are continuing to work to restore power at their data centre. Estimated completion time is 13:00 today. They have said that ADSL, Bonded DSL, and FTTC products connectivity has been restored. If your connection is still down please reboot your router. […]

Read More →